# Optimality of quasi-Monte Carlo methods and suboptimality of the sparse-grid Gauss–Hermite rule

### v1

@@ 4.1 Affinely-mapped QMC
$Q_{N}(f)=\frac{(2b)^{d}}{N}\sum_{j=1}^{N}f\bigl(\boldsymbol{\Psi}_{\mathrm{affine}}(\boldsymbol{t}^{(j)})\bigr)\rho\bigl(\boldsymbol{\Psi}_{\mathrm{affine}}(\boldsymbol{t}^{(j)})\bigr),\quad f\in H_{\rho}^{\alpha}.$ (10)
@@
Let $\alpha\in\mathbb{N}$, and let $Q_{N}$ be the rank-$1$ lattice rule defined in (10) with $b=\sqrt{\frac{\alpha}{2}\ln N}$. Then, with a generating vector $\boldsymbol{z}$ obtained by the CBC construction in [12, Algorithm 3.14], $Q_{N}$ satisfies
$\sup_{f\neq 0}\frac{|I(f)-Q_{N}(f)|}{\|f\|_{H_{\rho}^{\alpha}}}\leq C_{\alpha,d}N^{-\alpha}(\ln N)^{\frac{3\alpha d}{2}+\frac{d}{4}}.$ (11)
Proof.
First, we will show that $f\in H_{\rho}^{\alpha}$ implies $\|f\rho\|_{\text{unanchored}}+\|f\rho\|_{\text{decay}}\leq C_{d,\alpha}\|f\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}$ for some constant $C_{d,\alpha}>0$, where
$\|f\rho\|_{\text{decay}}\coloneqq\sup_{\begin{subarray}{c}\boldsymbol{x}\in\mathbb{R}^{d}\\ \mathbf{r}\in\{0,\ldots,\alpha-1\}^{d}\end{subarray}}\left|\rho^{-\varepsilon}(\boldsymbol{x})\,D^{\mathbf{r}}(f\rho)(\boldsymbol{x})\right|<\infty$
for all $\varepsilon\in(0,1/2)$, and
@@
$\langle f,h\rangle_{\mathscr{H}^{\alpha}([\boldsymbol{a},\boldsymbol{b}])}$ $:=\sum_{\begin{subarray}{c}\mathbf{r}\in\{0,\dots,\alpha\}^{d}\\ \mathfrak{v}:=\{k:r_{k}=\alpha\}\end{subarray}}\int_{[\boldsymbol{a}_{\mathfrak{v}},\boldsymbol{b}_{\mathfrak{v}}]}\left(\int_{[\boldsymbol{a}_{-\mathfrak{v}},\boldsymbol{b}_{-\mathfrak{v}}]}D^{\mathbf{r}}f(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}_{-\mathfrak{v}}\right)$ (12)
@@
for bounded $\boldsymbol{a},\boldsymbol{b}]$. In the proof we use the following identities
@@
$D^{\mathbf{r}}g$ $=\sum_{\mathbf{s}\leq\mathbf{r}}\binom{\mathbf{r}}{\mathbf{s}}(D^{\mathbf{s}}\rho)\,(D^{\mathbf{r}-\mathbf{s}}f)=\sum_{\mathbf{s}\leq\mathbf{r}}(-1)^{|\mathbf{s}|}\sqrt{\mathbf{s}!}\binom{\mathbf{r}}{\mathbf{s}}\,H_{\mathbf{s}}\,\rho\,D^{\mathbf{r}-\mathbf{s}}f.$ (13)
@@
$\sup_{[\boldsymbol{a},\boldsymbol{b}]\subset\mathbb{R}^{d}}\|g\|_{\mathscr{H}^{\alpha}([\boldsymbol{a},\boldsymbol{b}])}^{2}\leq\sum_{\begin{subarray}{c}\mathbf{r}\in\{0,\ldots,\alpha\}^{d}\\ \mathfrak{v}:=\left\{j:r_{j}=\alpha\right\}\end{subarray}}\int_{\mathbb{R}^{|\mathfrak{v}|}}\left(\int_{\mathbb{R}^{d-|\mathfrak{\mathfrak{v}}|}}|D^{\mathbf{r}}g(\boldsymbol{x})|\,\mathrm{d}\boldsymbol{x}_{-\mathfrak{v}}\right)^{2}\mathrm{d}\boldsymbol{x}_{\mathfrak{v}}.$ (14)
@@
with $\|H_{\boldsymbol{s}(-\mathfrak{v})}\|_{L^{2}_{\rho_{\mathfrak{-\mathfrak{v}}}}}:=\left(\int_{\mathbb{R}^{d-|\mathfrak{\mathfrak{v}}|}}|H_{\boldsymbol{s}(-\mathfrak{v})}(\boldsymbol{x}_{-\mathfrak{v}})|^{2}\,\rho_{\mathfrak{-\mathfrak{v}}}(\boldsymbol{x}_{-\mathfrak{v}})\mathrm{d}\boldsymbol{x}_{-\mathfrak{v}}\right)^{1/2}$. Here, we used the notations $H_{\boldsymbol{s}(\mathfrak{v})}(x_{\mathfrak{v}}):=\prod_{j\in\mathfrak{v}}H_{s_{j}}(x_{j})$, $H_{\boldsymbol{s}(-\mathfrak{v})}(x_{-\mathfrak{v}}):=\prod_{j\in\{1,\dots,d\}\setminus\mathfrak{v}}H_{s_{j}}(x_{j})$, and similarly $\rho_{\mathfrak{v}}:=\prod_{j\in\mathfrak{v}}\rho(x_{j})$ and $\rho_{-\mathfrak{v}}:=\prod_{j\in\{1,\dots,d\}\setminus\mathfrak{v}}\rho(x_{j})$. Noting $\|H_{\boldsymbol{s}(-\mathfrak{v})}\|_{L^{2}_{\rho_{\mathfrak{-\mathfrak{v}}}}}=1$, it follows
@@
$\leq C_{d}\sum_{\mathbf{s}\leq\mathbf{r}}\mathbf{s}!\binom{\mathbf{r}}{\mathbf{s}}^{2}\int_{\mathbb{R}^{d}}|H_{\boldsymbol{s}(\mathfrak{v})}(\boldsymbol{x}_{\mathfrak{v}})|^{2}\rho_{\mathfrak{\mathfrak{v}}}^{2}(\boldsymbol{x}_{\mathfrak{v}})\,|D^{\mathbf{r}-\mathbf{s}}f(\boldsymbol{x})|^{2}\,\rho_{\mathfrak{-\mathfrak{v}}}(\boldsymbol{x}_{-\mathfrak{v}})\mathrm{d}\boldsymbol{x}$
$\leq C_{d}\sum_{\mathbf{s}\leq\mathbf{r}}\mathbf{s}!\binom{\mathbf{r}}{\mathbf{s}}^{2}\sup_{\boldsymbol{y}_{\mathfrak{v}}\in\mathbb{R}^{|\mathfrak{v}|}}|H_{\boldsymbol{s}(\mathfrak{v})}(\boldsymbol{y}_{\mathfrak{v}})|^{2}\rho_{\mathfrak{\mathfrak{v}}}(\boldsymbol{y}_{\mathfrak{v}})\int_{\mathbb{R}^{d}}|D^{\mathbf{r}-\mathbf{s}}f(\boldsymbol{x})|^{2}\,\rho(\boldsymbol{x})\mathrm{d}\boldsymbol{x}.$
@@
$\sup_{[\boldsymbol{a},\boldsymbol{b}]\subset\mathbb{R}^{d}}\|g\|_{\mathscr{H}^{\alpha}([\boldsymbol{a},\boldsymbol{b}])}\leq C_{d,\alpha}\|f\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}$
@@
To show $\lVert g\rVert_{\text{decay}}\leq C\|f\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}$, we will invoke the Sobolev embedding in Proposition 2.3 to bound the supremum norm. Define $h_{\mathbf{r}}:=\rho^{-\varepsilon}D^{\mathbf{r}}g=\rho^{-\varepsilon}D^{\mathbf{r}}(f\rho)$ for $\mathbf{r}\in\{0,\dots,\alpha-1\}^{d}$, for which we will derive a bound on its Sobolev norm. For $\mathbf{k}\in\{0,1\}^{d}$
@@
$\qquad\qquad\times\|D^{\mathbf{r}+\mathbf{k}-\mathbf{j}-\mathbf{s}}f\|_{L_{\rho}^{2}(\mathbb{R}^{d})}<\infty,$ (15)
@@
$\sum_{\mathbf{k}\in\{0,1\}^{d}}\|D^{\mathbf{k}}(\rho^{-\varepsilon}D^{\mathbf{r}}g)\|_{L^{2}(\mathbb{R}^{d})}^{2}=\sum_{\mathbf{k}\in\{0,1\}^{d}}\|D^{\mathbf{k}}h_{\mathbf{r}}\|_{L^{2}(\mathbb{R}^{d})}^{2}$ $\leq C\|f\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}^{2}.$
Therefore, in view of Proposition 2.3 we have $\sup_{\boldsymbol{x}\in\mathbb{R}^{d}}|\rho^{-\varepsilon}(\boldsymbol{x})D^{\mathbf{r}}g(\boldsymbol{x})|\leq C\|f\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}$, hence $\|f\rho\|_{\text{unanchored}}+\|f\rho\|_{\text{decay}}\leq C_{d,\alpha}\|f\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}$ follows.
With this, from the first displayed equation in [55, Lemma 5] together with [55, Propositions 7 & 8], if $\boldsymbol{z}$ satisfies the second displayed equation in [55, Lemma 5], then this $\boldsymbol{z}$ satisfies (11); see the proof of [55, Theorem 2]. Since such $\boldsymbol{z}$ can be constructed using a CBC construction in [12, Algorithm 3.14], the proof is complete. ∎
Affinely mapped digital nets also achieve the optimal convergence rate, up to a logarithmic factor, as shown in [11, Corollary 1].
Proposition 4.2 (Affinely-mapped digital net, [11, Corollary 1]).
@@
In the result quoted above, higher-order nets of order $(2\alpha+1)$ are considered. On the other hand, as noted in [11], higher-order nets of order $\alpha$ also achieve the optimal polynomial rate $N^{-\alpha}$, but with a logarithmic factor of a larger exponent.
@@ 4.2 QMC with Möbius transformation
$\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t})\coloneqq(\phi(t_{1}),\ldots,\phi(t_{d})),\;\;\;\phi(t)\coloneqq-\cot(\pi t).$ (16)
This is a specific Möbius transformation considered in [64]. For the univariate transformation $\phi$, its derivatives satisfy the following.
@@
Proof.
Noting $\phi^{(\tau)}(t)=-\pi^{\tau}\cot^{(\tau)}(\pi t)$ and $\phi^{-1}(x)=\frac{1}{\pi}\cot^{-1}(-x)$, it suffices to show that $\cot^{(\tau)}\circ\cot^{-1}(x)$ is a polynomial of degree $\tau+1$. From [34, Equation (5)] we know that $P_{\tau}(x):=\tan^{(\tau)}\circ\tan^{-1}(x)$ is a polynomial of degree $\tau+1$ satisfying
@@
with $P_{0}(x)=x$ and $P_{1}(x)=x^{2}+1$. With $\cot(t)=\tan(\pi/2-t)$, the statement follows. ∎
Using $\boldsymbol{\Psi}_{\mathrm{cot}}$ in (9) yields the following formula:
$Q_{N}(f)=\frac{1}{N}\sum_{j=1}^{N}f\left(\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t}^{(j)})\right)\rho\left(\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t}^{(j)})\right)\prod_{k=1}^{d}\phi^{\prime}(t^{(j)}_{k}),\quad f\in H_{\rho}^{\alpha}.$ (17)
We derive a higher-order chain rule for component-wise composition, which may be of interest by itself. Although this is a special case of the general Faa di Bruno formula, component-wise composition makes the formula simpler and thus easier to work with for broader context in numerical analysis.
First, we derive a recurrence relation in the univariate setting. For $r\geq 1$ and $1\leq k\leq r$, let $B_{r,k}$ denote the (partial) Bell polynomial. We use the convention $B_{0,0}=1$ and $B_{r,0}=0$ for $r\geq 1$.
Define $C_{0,0}:=B_{0,0}=1$ and $C_{r,-1}:=0$ for all $r\in\mathbb{N}_{0}$. For $r\in\mathbb{N}$, $C_{r,0}:=B_{r,0}=0$, and for integers $1\leq\eta\leq r$, define
@@
Proof.
The cases $r=0$ as well as $r\in\mathbb{N}$ and $\eta=0$ are easy to check. We show the statement for $1\leq\eta\leq r$. First, the partial Bell polynomials satisfy the recurrence relation

### v2

@@ 4.1 Affinely-mapped QMC
(10) $Q_{N}(f)=\frac{(2b)^{d}}{N}\sum_{j=1}^{N}f\bigl(\boldsymbol{\Psi}_{\mathrm{affine}}(\boldsymbol{t}^{(j)})\bigr)\rho\bigl(\boldsymbol{\Psi}_{\mathrm{affine}}(\boldsymbol{t}^{(j)})\bigr),\quad f\in H_{\rho}^{\alpha}.$
@@
Let $\alpha\in\mathbb{N}$, and let $Q_{N}$ be the rank-$1$ lattice rule defined in (10) with $b={(2+\eta)}\sqrt{\alpha\ln N}$, where $\eta>0$ is arbitrary. Then, with a generating vector $\boldsymbol{z}$ obtained by the CBC construction in [undefl, Algorithm 3.14], $Q_{N}$ satisfies
(11) $\sup_{f\neq 0}\frac{|I(f)-Q_{N}(f)|}{\|f\|_{H_{\rho}^{\alpha}}}\leq C_{\alpha,d,{\eta}}\;\!N^{-\alpha}(\ln N)^{\frac{3\alpha d}{2}+\frac{d}{4}}.$
Proof 4.3.
First, we will show that $f\in H_{\rho}^{\alpha}$ implies $\|f\rho\|_{\text{unanchored}}+\|f\rho\|_{\text{decay}{,\varepsilon}}\leq C_{d,\alpha,{\varepsilon}}\|f\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}$ for every $\varepsilon\in(0,1/2)$ with some constant $C_{d,\alpha{,\varepsilon}}>0$, where
$\|f\rho\|_{\text{decay}{,\varepsilon}}\coloneqq\sup_{\begin{subarray}{c}\boldsymbol{x}\in\mathbb{R}^{d}\\ \mathbf{r}\in\{0,\ldots,\alpha-1\}^{d}\end{subarray}}\left|\rho^{-\varepsilon}(\boldsymbol{x})\,D^{\mathbf{r}}(f\rho)(\boldsymbol{x})\right|<\infty,$
and
@@
(12) $\langle f,h\rangle_{\mathscr{H}^{\alpha}([\boldsymbol{a},\boldsymbol{b}])}$ $:=\sum_{\begin{subarray}{c}\mathbf{r}\in\{0,\dots,\alpha\}^{d}\\ \mathfrak{v}:=\{k:r_{k}=\alpha\}\end{subarray}}\int_{[\boldsymbol{a}_{\mathfrak{v}},\boldsymbol{b}_{\mathfrak{v}}]}\left(\int_{[\boldsymbol{a}_{-\mathfrak{v}},\boldsymbol{b}_{-\mathfrak{v}}]}D^{\mathbf{r}}f(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}_{-\mathfrak{v}}\right)$
@@
for bounded $\boldsymbol{a},\boldsymbol{b}]$. The condition $\|f\rho\|_{\text{decay},\varepsilon}<\infty$ corresponds to the exponential decay condition [undefaae, Equation (22)], whereas $\|f\rho\|_{\text{unanchored}}<\infty$ corresponds to the requirement that [undefaae, Equation (24)] be finite. In the proof we use the following identities
@@
(13) $D^{\mathbf{r}}g$ $=\sum_{\mathbf{s}\leq\mathbf{r}}\binom{\mathbf{r}}{\mathbf{s}}(D^{\mathbf{s}}\rho)\,(D^{\mathbf{r}-\mathbf{s}}f)=\sum_{\mathbf{s}\leq\mathbf{r}}(-1)^{|\mathbf{s}|}\sqrt{\mathbf{s}!}\binom{\mathbf{r}}{\mathbf{s}}\,H_{\mathbf{s}}\,\rho\,D^{\mathbf{r}-\mathbf{s}}f.$
@@
(14) $\sup_{[\boldsymbol{a},\boldsymbol{b}]\subset\mathbb{R}^{d}}\|g\|_{\mathscr{H}^{\alpha}([\boldsymbol{a},\boldsymbol{b}])}^{2}\leq\sum_{\begin{subarray}{c}\mathbf{r}\in\{0,\ldots,\alpha\}^{d}\\ \mathfrak{v}:=\left\{j:r_{j}=\alpha\right\}\end{subarray}}\int_{\mathbb{R}^{|\mathfrak{v}|}}\left(\int_{\mathbb{R}^{d-|\mathfrak{\mathfrak{v}}|}}|D^{\mathbf{r}}g(\boldsymbol{x})|\,\mathrm{d}\boldsymbol{x}_{-\mathfrak{v}}\right)^{2}\mathrm{d}\boldsymbol{x}_{\mathfrak{v}}.$
@@
with $\|H_{\boldsymbol{s}(-\mathfrak{v})}\|_{L^{2}_{\rho_{\mathfrak{-\mathfrak{v}}}}}:=\left(\int_{\mathbb{R}^{d-|\mathfrak{\mathfrak{v}}|}}|H_{\boldsymbol{s}(-\mathfrak{v})}(\boldsymbol{x}_{-\mathfrak{v}})|^{2}\,\rho_{\mathfrak{-\mathfrak{v}}}(\boldsymbol{x}_{-\mathfrak{v}})\mathrm{d}\boldsymbol{x}_{-\mathfrak{v}}\right)^{1/2}$. Here, we used the notations $H_{\boldsymbol{s}(\mathfrak{v})}({\boldsymbol{x}}_{\mathfrak{v}}):=\prod_{j\in\mathfrak{v}}H_{s_{j}}(x_{j})$, $H_{\boldsymbol{s}(-\mathfrak{v})}({\boldsymbol{x}}_{-\mathfrak{v}}):=\prod_{j\in\{1,\dots,d\}\setminus\mathfrak{v}}H_{s_{j}}(x_{j})$, and similarly $\rho_{\mathfrak{v}}{(\boldsymbol{x_{\mathfrak{v}}})}:=\prod_{j\in\mathfrak{v}}\rho(x_{j})$ and $\rho_{-\mathfrak{v}}{(\boldsymbol{x_{-\mathfrak{v}}})}:=\prod_{j\in\{1,\dots,d\}\setminus\mathfrak{v}}\rho(x_{j})$. Noting $\|H_{\boldsymbol{s}(-\mathfrak{v})}\|_{L^{2}_{\rho_{\mathfrak{-\mathfrak{v}}}}}=1$, it follows
@@
$\leq C_{d,{\alpha}}\sum_{\mathbf{s}\leq\mathbf{r}}\mathbf{s}!\binom{\mathbf{r}}{\mathbf{s}}^{2}\int_{\mathbb{R}^{d}}|H_{\boldsymbol{s}(\mathfrak{v})}(\boldsymbol{x}_{\mathfrak{v}})|^{2}\rho_{\mathfrak{\mathfrak{v}}}^{2}(\boldsymbol{x}_{\mathfrak{v}})\,|D^{\mathbf{r}-\mathbf{s}}f(\boldsymbol{x})|^{2}\,\rho_{\mathfrak{-\mathfrak{v}}}(\boldsymbol{x}_{-\mathfrak{v}})\mathrm{d}\boldsymbol{x}$
$\leq C_{d,{\alpha}}\sum_{\mathbf{s}\leq\mathbf{r}}\mathbf{s}!\binom{\mathbf{r}}{\mathbf{s}}^{2}\sup_{\boldsymbol{y}_{\mathfrak{v}}\in\mathbb{R}^{|\mathfrak{v}|}}|H_{\boldsymbol{s}(\mathfrak{v})}(\boldsymbol{y}_{\mathfrak{v}})|^{2}\rho_{\mathfrak{\mathfrak{v}}}(\boldsymbol{y}_{\mathfrak{v}})\int_{\mathbb{R}^{d}}|D^{\mathbf{r}-\mathbf{s}}f(\boldsymbol{x})|^{2}\,\rho(\boldsymbol{x})\mathrm{d}\boldsymbol{x}.$
@@
${\|f\rho\|_{\text{unanchored}}=}\sup_{[\boldsymbol{a},\boldsymbol{b}]\subset\mathbb{R}^{d}}\|g\|_{\mathscr{H}^{\alpha}([\boldsymbol{a},\boldsymbol{b}])}\leq C_{d,\alpha}\|f\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}$
@@
To show $\lVert g\rVert_{\text{decay}{,\varepsilon}}\leq C\|f\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}$, we will invoke the Sobolev embedding in Proposition 2.4 to bound the supremum norm. Define $h_{\mathbf{r}}:=\rho^{-\varepsilon}D^{\mathbf{r}}g=\rho^{-\varepsilon}D^{\mathbf{r}}(f\rho)$ for $\mathbf{r}\in\{0,\dots,\alpha-1\}^{d}$, for which we will derive a bound on its Sobolev norm. For $\mathbf{k}\in\{0,1\}^{d}$
@@
(15) $\qquad\qquad\times\|D^{\mathbf{r}+\mathbf{k}-\mathbf{j}-\mathbf{s}}f\|_{L_{\rho}^{2}(\mathbb{R}^{d})}<\infty,$
@@
$\sum_{\mathbf{k}\in\{0,1\}^{d}}\|D^{\mathbf{k}}(\rho^{-\varepsilon}D^{\mathbf{r}}g)\|_{L^{2}(\mathbb{R}^{d})}^{2}=\sum_{\mathbf{k}\in\{0,1\}^{d}}\|D^{\mathbf{k}}h_{\mathbf{r}}\|_{L^{2}(\mathbb{R}^{d})}^{2}$ $\leq C_{{\varepsilon}}\|f\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}^{2}.$
Therefore, in view of Proposition 2.4 we have $\sup_{\boldsymbol{x}\in\mathbb{R}^{d}}|\rho^{-\varepsilon}(\boldsymbol{x})D^{\mathbf{r}}g(\boldsymbol{x})|\leq C_{{\varepsilon}}\|f\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}$, hence $\|f\rho\|_{\text{unanchored}}+\|f\rho\|_{\text{decay}{,\varepsilon}}\leq C_{d,\alpha,{\varepsilon}}\|f\|_{H_{\rho}^{\alpha}(\mathbb{R}^{d})}$ follows.
With this, from the first displayed equation in [undefaae, Lemma 5] together with [undefaae, Propositions 7 & 8], if $\boldsymbol{z}$ satisfies the second displayed equation in [undefaae, Lemma 5], then this $\boldsymbol{z}$ satisfies (11); see the proof of [undefaae, Theorem 2]. Since such $\boldsymbol{z}$ can be constructed using a CBC construction in [undefl, Algorithm 3.14], we have
$|I(f)-Q_{N}(f)|\leq C_{\alpha,d,\beta}\|f\|_{H^{\alpha}_{\rho}(\mathbb{R}^{d})}\bigg(\frac{(\ln N)^{\alpha d}}{N^{\alpha}}b^{d(\alpha+1/2)}+b^{d\alpha}\mathrm{e}^{-\beta b^{2}}+b^{d-2}\mathrm{e}^{-\beta b^{2}}\bigg),$
where $\beta:=\varepsilon/2\in(0,1/4)$ corresponds to $\beta$ in [undefaae], and the constant $C_{\alpha,d,\beta}$ is independent of $N$ and $b$. For $\eta>0$ arbitrary, we now choose $b=(2+\eta)\sqrt{\alpha\ln(N)}$. Then, since the bound above holds for any $1/(2+\eta)^{2}\leq\beta<1/4$, for such $\beta$ this choice of $b$ ensures $\mathrm{e}^{-\beta b^{2}}=N^{-\beta(2+\eta)^{2}\alpha}\leq N^{-\alpha}$. This completes the proof.
Affinely mapped digital nets also achieve the optimal convergence rate, up to a logarithmic factor, as shown in [undefk, Corollary 1].
Proposition 4.4 (Affinely-mapped digital net, [undefk, Corollary 1]).
@@
In the result quoted above, higher-order nets of order $(2\alpha+1)$ are considered. On the other hand, as noted in [undefk], higher-order nets of order $\alpha$ also achieve the optimal polynomial rate $N^{-\alpha}$, but with a logarithmic factor of a larger exponent.
@@ 4.2 QMC with Möbius transformation
(16) $\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t})\coloneqq(\phi(t_{1}),\ldots,\phi(t_{d})),\;\;\;\phi(t)\coloneqq-\cot(\pi t).$
This is a specific Möbius transformation considered in [undefaan]. Using $\boldsymbol{\Psi}_{\mathrm{cot}}$ in (9) yields the following formula:
(17) $Q_{N}(f)=\frac{1}{N}\sum_{j=1}^{N}f\left(\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t}^{(j)})\right)\rho\left(\boldsymbol{\Psi}_{\mathrm{cot}}(\boldsymbol{t}^{(j)})\right)\prod_{k=1}^{d}\phi^{\prime}(t^{(j)}_{k}),\quad f\in H_{\rho}^{\alpha}.$
It turns out that the transformed integrand, defined on the unit cube by $g:=f(\boldsymbol{\Psi}_{\mathrm{cot}}(\cdot))\rho(\boldsymbol{\Psi}_{\mathrm{cot}}(\cdot))\prod_{k=1}^{d}\phi^{\prime}$, belongs to suitable Sobolev spaces for which QMC methods work well. To show this, we first derive results on the derivatives of $g$.
The derivatives of the transformation $\phi(t)=-\cot(\pi t)$ satisfy the following.
@@
Proof 4.6.
Noting $\phi^{(\tau)}(t)=-\pi^{\tau}\cot^{(\tau)}(\pi t)$ and $\phi^{-1}(x)=\frac{1}{\pi}\cot^{-1}(-x)$, it suffices to show that $\cot^{(\tau)}\circ\cot^{-1}(x)$ is a polynomial of degree $\tau+1$. From [undefah, Equation (5)] we know that $P_{\tau}(x):=\tan^{(\tau)}\circ\tan^{-1}(x)$ is a polynomial of degree $\tau+1$ satisfying
@@
with $P_{0}(x)=x$ and $P_{1}(x)=x^{2}+1$. With $\cot(t)=\tan(\pi/2-t)$, the statement follows.
Next, we derive a higher-order chain rule for component-wise composition, which may be of interest by itself. Although this is a special case of the general Faa di Bruno formula, component-wise composition makes the formula simpler and thus easier to work with for broader context in numerical analysis.
First, we derive a recurrence relation in the univariate setting. For $r\geq 1$ and $1\leq{\eta}\leq r$, let $B_{r,{\eta}}$ denote the (partial) Bell polynomial, which may be defined by
(18) $B_{r,\eta}(x_{1},\dots,x_{r})=\sum_{\boldsymbol{0}\leq\mathbf{k}\leq\boldsymbol{\eta}}r!\bigg(\prod_{\ell=1}^{r}\frac{x_{\ell}^{k_{\ell}}}{k_{\ell}!(\ell!)^{k_{\ell}}}\bigg)\mathbbm{1}_{\{|\mathbf{k}|=\eta,\,\sum_{\ell=1}^{r}\ell k_{\ell}=r\}}(\mathbf{k}),$
where $\mathbbm{1}_{A}(\mathbf{k})$ denotes the indicator function of a set $A$, and $\boldsymbol{\eta}:=(\eta,\ldots,\eta)$; see for example [undeff, Definition 11.2]. We use the convention $B_{0,0}=1$ and $B_{r,0}=0$ for $r\geq 1$. We note that although $B_{r,\eta}=B_{r,\eta}(x_{1},\dots,x_{r})$ is defined as a polynomial in $x_{1},\dots,x_{r}$, the constraints $\sum_{i=1}^{r}k_{i}=\eta$ and $\sum_{i=1}^{r}ik_{i}=r$ in its definition imply that $B_{r,\eta}$ in fact depends only on $x_{1},\dots,x_{r-\eta+1}$.
We will use the following identity, which shows that $\partial_{j}B_{r,\eta}$ is, up to a binomial coefficient, again a partial Bell polynomial. Although the result appears to be well known, a proof seems difficult to find in the literature, so we provide one.
Lemma 4.7.
Let $r\in\mathbb{N}$ and $1\leq\eta\leq r$. Then, for $j=1,\dots,r-\eta+1$, we have
$\partial_{j}B_{r,\eta}(x_{1},\ldots,x_{r-\eta+1})=\binom{r}{j}\,B_{r-j,\eta-1}\!\left(x_{1},\ldots,x_{r-j-(\eta-1)+1}\right).$
Proof 4.8.
If we differentiate $B_{r,\eta}$ defined in (18) with respect to $x_{j}$, the terms with $k_{j}=0$ vanish: $\partial_{j}B_{r,\eta}=\sum_{\mathbf{e}_{j}\leq\mathbf{k}\leq\boldsymbol{\eta}}r!\frac{k_{j}\,x_{j}^{k_{j}-1}}{k_{j}!(j!)^{k_{j}}}\big(\prod_{\begin{subarray}{c}\ell=1\\ \ell\neq j\end{subarray}}^{r}\frac{x_{\ell}^{k_{\ell}}}{k_{\ell}!(\ell!)^{k_{\ell}}}\big)\mathbbm{1}_{\{|\mathbf{k}|=\eta,\,\sum_{\ell=1}^{r}\ell k_{\ell}=r\}}(\mathbf{k})$, where $\mathbf{e}_{j}$ is the unit vector whose $j$-th component is $1$ and whose other components are 0. We then extract the binomial coefficient $\binom{r}{j}=\frac{r!}{j!(r-j)!}$ as follows:
$\partial_{j}B_{r,\eta}$ $=\sum_{\mathbf{e}_{j}\leq\mathbf{k}\leq\boldsymbol{\eta}}\binom{r}{j}\frac{(r-j)!x_{j}^{k_{j}-1}}{(k_{j}-1)!(j!)^{k_{j}-1}}\bigg(\prod_{\begin{subarray}{c}\ell=1\\ \ell\neq j,\end{subarray}}^{r}\frac{x_{\ell}^{k_{\ell}}}{k_{\ell}!(\ell!)^{k_{\ell}}}\bigg)\mathbbm{1}_{\{|\mathbf{k}|=\eta,\,\sum_{\ell=1}^{r}\ell k_{\ell}=r\}}(\mathbf{k}).$
Re-indexing $k_{j}$ to $k_{j}+1$ yields
$\partial_{j}B_{r,\eta}$ $=\binom{r}{j}\sum_{\mathbf{0}\leq\mathbf{k}\leq\boldsymbol{\eta}-\mathbf{e}_{j}}(r-j)!\frac{x_{j}^{k_{j}}}{k_{j}!(j!)^{k_{j}}}\bigg(\prod_{\begin{subarray}{c}\ell=1\\ \ell\neq j\end{subarray}}^{r}\frac{x_{\ell}^{k_{\ell}}}{k_{\ell}!(\ell!)^{k_{\ell}}}\bigg)\mathbbm{1}_{\{|\mathbf{k}|=\eta-1,\,\sum_{\ell=1}^{r}\ell k_{\ell}=r-j\}}(\mathbf{k})$
$=\binom{r}{j}\sum_{\mathbf{0}\leq\mathbf{k}\leq\boldsymbol{\eta}-\mathbf{1}}(r-j)!\frac{x_{j}^{k_{j}}}{k_{j}!(j!)^{k_{j}}}\bigg(\prod_{\begin{subarray}{c}\ell=1\\ \ell\neq j\end{subarray}}^{r}\frac{x_{\ell}^{k_{\ell}}}{k_{\ell}!(\ell!)^{k_{\ell}}}\bigg)\mathbbm{1}_{\{|\mathbf{k}|=\eta-1,\,\sum_{\ell=1}^{r}\ell k_{\ell}=r-j\}}(\mathbf{k})$
$=\binom{r}{j}B_{r-j,\eta-1},$
where in the second line we used the fact that given the constraint $|\mathbf{k}|=\eta-1$, the range of summation may be replaced by $\mathbf{0}\leq\mathbf{k}\leq\boldsymbol{\eta}-\mathbf{1}$. This completes the proof.
To derive a higher-order chain rule, we consider the functions $C_{r,\eta}(t)$, defined by evaluating partial Bell polynomials at the derivatives of $\Phi$. Define $C_{0,0}:=B_{0,0}=1$ and $C_{r,-1}:=0$ for all $r\in\mathbb{N}_{0}$. For $r\in\mathbb{N}$, $C_{r,0}:=B_{r,0}=0$, and for integers $1\leq\eta\leq r$, define
@@
Proof 4.10.
The cases $r=0$ as well as $r\in\mathbb{N}$ and $\eta=0$ are easy to check. We show the statement for $1\leq\eta\leq r$. First, the partial Bell polynomials satisfy the recurrence relation [undeff, Equation (11.11)],